6. A firetruck going at 40m/s with a 900Hz siren is approaching a car from behind. The car going 20m/s (Both firetruck and car are initially going in the same direction). What is the frequency experienced, fo, for the person in the car? a, 900 Hz b. 925 Hz ?. 853 Hz d. 959 Hz e. 1001 Hz 7. Using the exact same speeds as in Problem #6 above and the same siren frequency, what is the frequency experienced, fo, for the person in the car after the firetruck has passed? a. 900 Hz b. 925 Hz ?. 825 Hz d. 853 Hz e. 980 Hz
Answer #1

the speed of sound in air is 343 m/s

6) d) The frequency experienced by the person in the car when they are moving in the same direction and the truck is approaching the car is

f_0=900\frac{343-20}{343-40}=959\,Hz

7) d) When the truck has passed, the car is moving towards the truck and the truck is moving away from the car. So the signs of the velocity of car and truck in the above equation flips

f_0=900\frac{343+20}{343+40}=853\,Hz